# Aliasing and Label-Independent Decomposition of Risk: Beyond the bias–variance trade-off

### v1

@@ Abstract
A central problem in data science is to use potentially noisy samples of an unknown function to predict function values for unseen inputs. In classical statistics, the predictive error is understood as a trade-off between the bias and the variance that balances model simplicity with its ability to fit complex functions. However, over-parameterized models exhibit counter-intuitive behaviors, such as “double descent” in which models of increasing complexity exhibit decreasing generalization error. We introduce an alternative paradigm called the generalized aliasing decomposition. We explain the asymptotically small error of complex models as a systematic “de-aliasing” that occurs in the over-parameterized regime. In the limit of large models, the contribution due to aliasing vanishes, leaving an expression for the asymptotic total error we call the invertibility failure of very large models on few training points. Because the generalized aliasing decomposition can be explicitly calculated from the relationship between model class and samples without seeing any data labels, it can answer questions related to experimental design and model selection before collecting data or performing experiments. We demonstrate this approach using several examples, including classical regression problems and a cluster expansion model used in materials science.
@@ I Introduction
Simple models are generally preferred for many reasons, including interpretability and computational expense [1, 2, 3, 4, 5, 6, 7], but one of the more important justifications for parsimony is a need to balance over- and under-fitting. Models with few parameters avoid making wild predictions but coarsely interpolate the observed data without much fidelity (high bias), while over-parameterized models fit the sampled data well with wild swings in between data points (high variance). The unquestioned goal has been to find the “sweet spot” of model complexity that balances bias and variance, i.e., a faithful model of marginal complexity (see Figure 1, left panel).
@@
While the foregoing story has long been the central dogma of statistics, we now know this view of the fitting problem is not the whole story (see [8] for the bias-variance decomposition for neural networks and [9] for a more thorough presentation of the topic). For extremely over-parameterized models (i.e., more parameters than samples), prediction errors often actually decrease with additional parameters. This phenomenon was recently coined “double descent”[9], although aspects of the phenomena had been known much earlier [10]. More generally, prediction errors may exhibit complex, “multiple descents” [11, 12, 13, 14, 15, 16, 17, 18].
@@
Nonconvex risk curves are most famously recognized in neural networks[19, 20], though this behavior has been observed in various other settings as well (see [10] for a thorough review). Several recent attempts have been made to explain this behavior, focusing on regression and the simplest case of “double descent”, although recent studies show that risk curves may be far more complicated [11]. In [14, 17], the bias–variance decomposition is expanded to explain this non-convex behavior, relying on the interplay between the model design and the actual data. Several other efforts have been made to clarify the relationship between the model class, inherent algorithmic bias, testing and training data, and the appearance of nonmonotonic loss and generalization curves. We do not present a comprehensive summary, but direct the interested reader to [21, 13, 12, 18] that clarify the nature of double descent and its apparent reliance on the structure of the testing and training data sets.
@@
In contrast to these approaches, we build on insights from signal processing [20] and introduce a novel decomposition (34), which we refer to as the generalized aliasing decomposition summarized in the right panel of Figure 1.
Figure 1: (Left panel) Generalization errors are traditionally understood in terms of the bias–variance trade-off which are difficult to predict in the modern “interpolating” regime. (Right panel) The Generalized Aliasing Decomposition explains generalization curves as contributions from Aliasing (blue) and Invertibility (red). The former explains the structure of mid-sized models, while the latter explains the error for both very small and very large models. Aliasing is further decomposed as contributions coming from the (pseudo) inverse of the traditional design matrix (dashed blue) that account for nonmonotonic structure in the generalization curve and a monotonically decreasing nescience transform that accounts for the scale. The invertibility error is further decomposed in terms of a monotonically increasing parameter bias and monotonically deceasing contributions from nescient parameters. This intuitive, interpretable decomposition applies throughout both classical and modern regimes.
Using this decomposition, we demonstrate that a nonconvex generalization curve is a generic phenomenon, which includes, as a special case, aliasing in the Fourier representation of a time series wherein high-frequency (noisy) components of a signal alias with the lower-frequency (modeled) components, corrupting the representation. Aliasing errors can often be reduced by including more terms in the model class than are required simply from the dimensionality of the data.
The aliasing error is decomposed as the product of the inverse of the classical design matrix and a term we call the nescience transform. The contribution from the design matrix explains much of the nonmonotonic behavior of the generalization curve. Model error due to aliasing increases when the most recently added model terms are linearly dependent on already fitted aspects of the model in the sampled subspace. While this aliasing error is generally maximized at the interpolation threshold, leading to the simplest version of double descent, it is not necessarily maximized there, and real-world sampling strategies can produce complex but understandable error curves (see Figure 4, for example). The scale of the aliasing errors is determined by the norm of the nescience transform. Finally, the asymptotic error is determined by an invertibility term, which can be further decomposed as a sum of bias in the modeled parameters and contributions from the nescient parameters.
@@
Taken collectively, this decomposition explains all the qualitative features of generic risk curves. It further clarifies the roles of model structure, data sampling, data labels, and the learning algorithm in a way that is intuitive while formally facilitating key modeling decisions such as the choice of model class, experimental design, regularization, and learning algorithm.
I.1 Aliasing and Decomposing Risk
@@
As an illustration, consider a polynomial fit on an interval $[a,b]$, and take the basis functions to be the usual monomial basis $\{1,t,t^{2},t^{3},\dots,t^{d}\}$ for some $d>0$; so $\Phi_{j}=t^{j-1}$, and the design matrix $\mathbf{M}$ is the Vandermonde matrix [^1] Although fitting monomials is the canonical pedagogical example, the ill-conditioned Vandemonde matrix makes this basis ill-suited for practical applications.
@@
$\hat{f}(t)=\sum_{j}\Phi_{j}(t)\hat{\bm{\theta}}_{j}.$ (4)
A primary quantity of interest is the so-called generalization error or population risk
$R_{\bm{\theta}}(\hat{f})=\mathbb{E}_{t}[(y(t)-\hat{f}(t))^{2}],$ (5)
where the expectation $\mathbb{E}$ is taken over a (usually theoretical) distribution of all of the data (not just the training samples), and the dependence of $R_{\bm{\theta}}$ on the model class itself is implicit. A primary goal in data science is to identify the model class and degree of complexity that minimizes risk (5).
II Results
With the context given above in place, we now describe our general results. We first assume that the model functions $\Phi$ may be extended to form a complete set over the space of all possible predictions (including the training points). The Weierstrass approximation theorem [23] guarantees that any continuous function on a closed interval $[a,b]$ can be approximated arbitrarily well by a polynomial; so if the full signal (modeled and unmodeled) is a continuous function, then the infinite monomial basis $\{1,t,t^{2},\dots\}$ is complete, and an appropriate extension for our example.
Because the basis functions $\Phi$ are complete, we can also represent the noise component $\xi$ of the signal as an expansion in $\Phi$. Thus, the fundamental ansatz for our analysis is
@@ II Results
We use the word nescience in the definition to emphasize the fact that the unmodeled part of the function $f$ is effectively unknown to us (noise, for example). Because $\mathbf{M}$ is bounded and linear, the subblocks $\mathbf{M}_{\mathcal{T}\mathcal{U}}$, $\mathbf{M}_{\mathcal{P}\mathcal{M}}$, and $\mathbf{M}_{\mathcal{P}\mathcal{U}}$ are also bounded linear transformations.
@@
In the case of Fourier series, the concept of aliasing refers to the distortion of a low-frequency signal by high-frequency modes. Expressed in the form we have described, Fourier aliasing is found from $\mathbf{A}=\mathbf{M}_{\mathcal{T}\mathcal{M}}^{+}\mathbf{M}_{\mathcal{T}\mathcal{U}}$, expressed in the Fourier basis for uniform samples (see Section II.II.3 for more on aliasing for Fourier series), where it can be expressed in closed-form. More generally, $\mathbf{A}$ quantifies how unmodeled components affect the signal at the sampled points, leading to a misrepresentation of the inferred modeled parameters that we call generalized aliasing.
II.1 Error Decomposition and Analysis
The model described above uses $\hat{\bm{\theta}}=(\hat{\bm{\theta}}_{\mathcal{M}},\bm{0})$ to approximate the true function $\bm{y}$ with the model output:
@@ II.1 Error Decomposition and Analysis
$\mathbf{E}_{\bm{\theta}}=\left(\begin{array}[]{cc}\mathbf{I}_{m}-\mathbf{B}&-\mathbf{A}\\ 0&\mathbf{I}\end{array}\right).$ (30)
@@
II.1.1 Bias and Variance
For a fixed decomposition $\Theta=\mathcal{M}\oplus\mathcal{U}$, the estimator $\hat{\bm{y}}$ and the operators $\mathbf{A}$, $\mathbf{B}$, and $\mathbf{E}_{\bm{\theta}}$ depend on the choice $T=\{t_{1},\dots,t_{n}\}$ of the training points. The expected value $\mathbb{E}_{T}(R_{\bm{\theta}}(\hat{\bm{y}}))=\mathbb{E}_{T}\left[\|\hat{\bm{y}}-\bm{y}\|^{2}\right]$ of the risk (taken over the distribution of $T$) is often decomposed into a sum of a bias term and a variance term; see, for example, [29, §20.1]. In our current setting, this decomposition is not intuitive and is difficult to analyze. We find it much more natural, and easier, to analyze the risk $\|\bm{y}-\hat{\bm{y}}\|^{2}$ and its expected value $\mathbb{E}_{T}\!\left[\|\bm{y}-\hat{\bm{y}}\|^{2}\right]$ directly, rather than to analyze the bias and variance terms separately.
II.1.2 Aliasing and Invertibility Error
A central question is how the risk changes with the dimension $m$ of the modeled space $\mathcal{M}$. Observe that
@@ II.1.2 Aliasing and Invertibility Error
We assume the transformation $\mathbf{M}$ has a bounded norm and note that its norm is independent of the choice of model $\mathcal{M}$. The overall error certainly depends on $\bm{\theta}$ and its norm, but, as we show below, the non-convex generalization is largely driven by the parameter error operator $\mathbf{E}_{\bm{\theta}}$, rather than by specific choices of $\bm{\theta}$. We focus now on $\mathbf{E}_{\bm{\theta}}$ and write it as a sum of a term proportional to $\mathbf{A}$ and the remainder:
$\mathbf{E}_{\bm{\theta}}$ $=\left(\begin{array}[]{cc}0&-\mathbf{A}\\ 0&0\end{array}\right)+\left(\begin{array}[]{cc}I-\mathbf{B}&0\\ 0&I\end{array}\right)$ (34)
$=\mathbf{E}_{\mathbf{A}}+\mathbf{E}_{\mathbf{B}}$ (35)
The operator $\mathbf{E}_{\mathbf{A}}$ measures the effect of aliasing, and the operator $\mathbf{E}_{\mathbf{B}}$ essentially measures two things: first, the failure of $\mathbf{M}_{\mathcal{T}\mathcal{M}}$ to be invertible (in the upper-left block), and second, how much of the signal is nescient, i.e., unmodeled (the lower-right identity block). The general shape of these different contributions to the risk is illustrated in the right panel of Figure 1. We now analyze each of $\mathbf{E}_{\mathbf{A}}$ and $\mathbf{E}_{\mathbf{B}}$ in turn.
II.1.3 Analysis of $\|\mathbf{E}_{\mathbf{A}}\|$ and $\|\mathbf{E}_{\mathbf{A}}\,\bm{\theta}\|$
Assuming that $\bm{\theta}$ and $\mathbf{M}$ are fixed with bounded norms, the norm $\|\mathbf{M}_{\mathcal{T}\mathcal{U}}\|$ of the nescience matrix is bounded by the norm of $\mathbf{M}$ and is a nonincreasing function of the model dimension $m$ (see Section IV.IV.2.1 for details), which means that the structure of
$\|\mathbf{E}_{\mathbf{A}}\bm{\theta}\|$ $=\|\mathbf{A}\bm{\theta}_{\mathcal{U}}\|=\|\mathbf{M}_{\mathcal{T}\mathcal{M}}^{+}\mathbf{M}_{\mathcal{T}\mathcal{U}}\bm{\theta}_{\mathcal{U}}\|$
$\leq\|\mathbf{M}_{\mathcal{T}\mathcal{M}}^{+}\|\|\mathbf{M}_{\mathcal{T}\mathcal{U}}\|\|\bm{\theta}_{\mathcal{U}}\|$
as a function of $m$ is primarily determined by the norm of the inverse design matrix $\mathbf{M}_{\mathcal{T}\mathcal{M}}^{+}$.

### v2

@@ Abstract
A central problem in data science is to use potentially noisy samples of an unknown function to predict function values for unseen inputs. In classical statistics, the predictive error is understood as a trade-off between the bias and the variance that balances model simplicity with its ability to fit complex functions. However, over-parameterized models exhibit counter-intuitive behaviors, such as “double descent” in which models of increasing complexity exhibit decreasing generalization error. In contrast to the bias-variance trade-off, we introduce an alternative paradigm called the generalized aliasing decomposition (GAD). We explain the asymptotically small error of complex models as a systematic “de-aliasing” that occurs in the over-parameterized regime. In the limit of large models, the error contribution due to aliasing vanishes, leaving an expression for the asymptotic total error we call the data insufficiencyfailure of very large models on few training points. Because the generalized aliasing decomposition can be explicitly calculated from the relationship between model class and samples without seeing any data labels, it can answer questions related to experimental design and model selection before collecting data or performing experiments. We demonstrate this approach using several examples, including classical regression problems and a cluster expansion model used in materials science.
Double descent |||| Benign overfitting |||| Risk |||| bias–variance trade-off |||| Interpolation |||| Implicit bias
I Introduction
@@
Some of the fundamental choices when model building are (1) the sample data and (2) the complexity of the model class. Simple models are generally preferred for many reasons, including interpretability and computational expense [1, 2, 3, 4, 5, 6, 7], but one of the more pragmatic justifications for parsimony is a need to balance over- and under-fitting. Models with few parameters avoid making wild predictions but under-fit the observed data without much fidelity (high bias), while over-parameterized models fit the sampled data well with wild swings in between data points (high variance). The unquestioned goal has been to find the “sweet spot” of model complexity that balances bias and variance, i.e., a faithful model of marginal complexity (see Figure 1, left panel) that minimizes the so-called “risk” or errors made by model predictions on unseen data.
@@
While the foregoing story has long been the central dogma of statistics, we now know this view of the fitting problem is not the whole story. For extremely over-parameterized models (i.e., more parameters than samples), prediction errors may actually decrease with additional parameters, a phenomenon recently coined “double descent”[8], summarized by the left panel in Figure 1. The boundary between the two regimes, known as the “interpolation threshold”, occurs when the model can perfectly interpolate the training data, naively when there are as many parameters as data. Non-convex risk curves are most famously recognized in neural networks[9, 10], though this behavior has been observed in various other settings as well. (See [11] for the bias-variance decomposition for neural networks and [12] for a thorough review.) Furthermore, models and data sets can be designed to exhibit complex, “multiple descents” [13, 14, 15, 16, 17, 18, 19, 20].
Although the rise of machine learning and big data has raised awareness of this phenomenon, the basic aspects have been known for some time in diverse fields [12]. Furthermore, highly structured data from scientific and engineering applications often exhibit similarly counter-intuitive behavior. The risk curve in the right panel of Figure 1 comes from a cluster expansion model of the formation enthalpy of alloy structures (see section III.3 for more detail). Not only do the peaks and troughs appear to the left of the interpolation threshold where classical bias-variance arguments ought to apply, the naive interpolation threshold apparently plays no role.
Figure 1: Limitations of the Bias-Variance Trade-off. Left: Bias and variance are traditionally understood an monotonically decreasing/increasing contributions to risk to be balanced by tuning model complexity. Double descent illustrates a breakdown of this intuition beyond the interpolation threshold where variance and bias can have exhibit counter-intuitive dependence on model class. Right: Highly structured data such as a cluster expansion of alloy formation enthalpy exhibit even more complicated and counter-intuitive dependence on model complexity.
More broadly, traditional data analysis techniques are also at odds with the intuition of the bias-variance decomposition. The Discrete Fourier Transform (DFT), for example, is formally equivalent to a regression problem (see section III.2) with as many parameters (Fourier coefficients) as data, so bias-variance arguments suggest that the inferred Fourier coefficients should exhibit unreasonably sensitivity to noisy data. In spite of this, the Fast Fourier Transform as a solution algorithm for the DFT at the interpolation threshold is one of the most influential and widely used algorithms in all of science and engineering (even for noisy signals that are not band-limited). Furthermore, techniques such as pseudo-spectral and collocation methods for solving differential equations are similarly equivalent to regression problems that are known to exhibit optimal performance beyond the interpolation threshold.
While the bias-variance decomposition holds as a formal mathematical result, the foregoing examples nevertheless reveal the limited insight it imbues. Its utility derives from the (incorrect) expectation that model selection balances the trade-off between monotonically decreasing (bias) and increasing (variance) error contributions. In reality, however, the contributions of bias and variance for each of the preceding examples are non-monotonic, complex, and intimately connected with the algorithmic solution to the optimization problem.
@@
Several recent attempts have been made to explain these behaviors, often focusing on regression and the simplest case of “double descent”, although recent studies show that risk curves may be far more complicated [13]. In [16, 19], the bias–variance decomposition is expanded to explain this nonconvex behavior, relying on the interplay between the model design and the actual data. Several other efforts have been made to clarify the relationship between the model class, inherent algorithmic bias, testing and training data, and the appearance of nonmonotonic loss and generalization curves. We do not present a comprehensive summary, but direct the interested reader to [21, 15, 14, 20] that clarify the nature of double descent and its apparent reliance on the structure of the testing and training data sets.
@@
In contrast to these approaches, we build on insights from signal processing [10] and introduce a novel decomposition (30), which we refer to as the generalized aliasing decomposition (GAD) summarized for the generic case of double descent in the left panel of Figure 2. The aliasing decomposition explains generic risk curves throughout both the classical and modern regimes as the contribution of three terms: parameter insufficiency, data insufficiency, and generalized aliasing.
Figure 2: Generalized Aliasing Decomposition. Left: The generalized aliasing decomposition (GAD) expresses risk as the contribution of three terms: model insufficiency, data insufficiency, and generalized aliasing. Model insufficiency dominates for small models and decreases monotonically with the number of parameters. Data insufficiency dominates for large models, increasingly monotonically with the number of parameters. Generalized Aliasing accounts for non-convex intermediate behaviors, but has a single peak at the interpolation threshold for the generic case, accounting for the phenomenon of Double Descent. Right: For highly structured problems (see Figure 1, right), aliasing explains all of the non-convex behavior of generic risk curves at intermediate model sizes.
Parameter insufficiency quantifies the inability of the model to fit the data. It is the dominate error contribution for models with few parameters and decreases monotonically with the number of parameters. Though the mapping is not exact, it roughly corresponds to “bias” in the bias-variance paradigm. Intuitively, adding more parameters to a model does not limit the ability of the model to fit data, so it decreases monotonically as we prove in the sequel.
Data insufficiency quantifies how much model parameters are unconstrained by available data. It is the dominate error contribution for models with many parameters, increasing monotonically as the model grows. Intuitively, adding more parameters does not introduce any additional parametric constraints, and we show this contribution does not increase with additional parameters.
Finally, generalized aliasing explains all non-monotonic behavior in intermediate regimes. It quantifies the extent to which independent components of the full signal are indistinguishable from the sampled data. The name derives from the special case of Fourier aliasing, in which high-frequency (noisy) components of a signal cannot be distinguished from low-frequency components at finite sampling rates. Thus high-frequency (unmodeled) contributions are said to alias with the low-frequency (modeled) components, corrupting the representation. In the generic case, aliasing errors are maximized at the interpolation threshold but can be reduced by including more terms in the model class. This generic scenario is responsible for the phenomenon of double descent (Figure 2 left). However, in structured cases such as the cluster expansion example (section III.3), aliasing fully accounts for the complicated, non-monotonic behavior throughout both the classical and modern regimes (Figure 2, right).
Furthermore, as we demonstrate below, aliasing is the intuition behind best-practices for other analysis techniques, such as pseudo-spectral approaches to solving differential equations. Although generalized aliasing is non-monotonic in the number of parameters, we show its behavior is straightforward to intuit based on whether or not the new model terms are independent of the current model on the sampled data. In section III.3, we use this fact to explain the apparently complicated risk curve in the right panel of Figure 1.
@@
Taken collectively, the three components of the GAD explain all the qualitative features of generic risk curves. GAD further clarifies the roles of model structure, data sampling, data labels, and the learning algorithm in an intuitive way. Indeed, a useful feature of the decomposition is that much of the analysis can be done independently of data labels.Consequently, the GAD formally facilitates key modeling decisions such as the choice of model class, experimental design, regularization, and learning algorithm.
II The Generalized Aliasing Decomposition (GAD)
II.1 Mathematical Preliminaries
@@
As an illustration, consider a polynomial fit on an interval $[a,b]$, and take the basis functions to be the usual monomial basis $\{1,t,t^{2},t^{3},\dots,t^{d}\}$ for some $d>0$; so $\Phi_{j}=t^{j-1}$, and the design matrix $\mathbf{M}$ is [^1] Although fitting monomials is the canonical pedagogical example, the ill-conditioning of this Vandermonde matrix makes this basis ill-suited for practical applications.
@@
$\hat{y}(t)$ $=\sum_{j}\Phi_{j}(t)\hat{\bm{\theta}}_{j}.$ (4)
An important quantity of interest is the usual mean squared error averaged over a (usually theoretical) distribution of all of the data (not just the training samples). The expectation of this the so-called generalization error or population risk is
$R_{\bm{\theta}}(\hat{\bm{y}})=\mathbb{E}[(\bm{y}-\hat{\bm{y}})^{2}]$ (5)
where the dependence of $R_{\bm{\theta}}$ on the model class itself is implicit. A primary goal in data science is to identify the model class and degree of complexity that minimizes risk (5).
II.2 The Generalized Aliasing Operator
With a common vocabulary established, we now derive the aliasing operating that underpins the GAD. We first assume that the model functions $\Phi$ may be extended to form a complete set over the space of all possible predictions (including the training points)222 The Weierstrass approximation theorem [23] guarantees that any continuous function on a closed interval $[a,b]$ can be approximated arbitrarily well by a polynomial; so if the full signal (modeled and unmodeled) is a continuous function, then the infinite monomial basis $\{1,t,t^{2},\dots\}$ is complete, and an appropriate extension for our example. Because the basis functions $\Phi$ are complete, we can also represent the noise component $\xi$ of the signal as an expansion in $\Phi$. Thus, the fundamental ansatz for our analysis is
@@ II.2 The Generalized Aliasing Operator
We use the word nescience in the definition to emphasize the fact that the unmodeled part of the function $f$ is effectively unknown to us (noise, for example). We further emphasize that with this definition, $\mathbf{M}$ does not denote the classical design matrix. Rather, $\mathbf{M}$ conceptually represents a basis transformation on the complete signal (including both seen and unseen data). Because $\mathbf{M}$ is bounded and linear, the subblocks $\mathbf{M}_{\mathcal{T}\mathcal{U}}$, $\mathbf{M}_{\mathcal{P}\mathcal{M}}$, and $\mathbf{M}_{\mathcal{P}\mathcal{U}}$ are also bounded linear transformations.
@@
In the case of Fourier series, the concept of aliasing refers to the distortion of a low-frequency signal by high-frequency modes. Expressed in the form we have described, Fourier aliasing is found from $\mathbf{A}=\mathbf{M}_{\mathcal{T}\mathcal{M}}^{+}\mathbf{M}_{\mathcal{T}\mathcal{U}}$, expressed in the Fourier basis for uniform samples (see Section II.2.III.2 for more on aliasing for Fourier series), where it can be expressed in closed-form. More generally, $\mathbf{A}$ quantifies how unmodeled components affect the signal at the sampled points, leading to a misrepresentation of the inferred modeled parameters that we call generalized aliasing.
Using $\hat{\bm{\theta}}=(\hat{\bm{\theta}}_{\mathcal{M}},\bm{0})$, we reconstruct the signal over both training and prediction points
@@
$\mathbf{E}_{\bm{\theta}}=\left(\begin{array}[]{cc}\mathbf{I}_{m}-\mathbf{B}&-\mathbf{A}\\ 0&\mathbf{I}_{\mathcal{U}}\end{array}\right).$ (30)
@@
The block decomposition in Eq. (30) is the generalized aliasing decomposition. The three nonzero blocks correspond to data insufficiency $\mathbf{I}_{m}-\mathbf{B}$, parameter insufficiency $\mathbf{I}_{\mathcal{U}}$, and generalized aliasing $-\mathbf{A}$, respectively as first presented in the left panel of Figure 2. We now analyze each of these contributions to the error in turn.
II.3 Error Analysis
For a fixed decomposition $\Theta=\mathcal{M}\oplus\mathcal{U}$, the estimator $\hat{\bm{y}}$ and the operators $\mathbf{A}$, $\mathbf{B}$, and $\mathbf{E}_{\bm{\theta}}$ depend on the choice $T=\{t_{1},\dots,t_{n}\}$ of the training points. The expected value $\mathbb{E}_{T}(R_{\bm{\theta}}(\hat{\bm{y}}))=\mathbb{E}_{T}\left[\|\hat{\bm{y}}-\bm{y}\|^{2}\right]$ of the risk (taken over the distribution of $T$) is often decomposed into a sum of a bias term and a variance term; see, for example, [30, §20.1]. In the present context, this decomposition is not intuitive and is difficult to analyze. We find it much more natural, and easier, to analyze the risk $\|\bm{y}-\hat{\bm{y}}\|^{2}$ and its expected value $\mathbb{E}_{T}\!\left[\|\bm{y}-\hat{\bm{y}}\|^{2}\right]$ directly, rather than to analyze the bias and variance terms separately.
We are primarily interested in estimating the risk for a particular decomposition. Observe that
@@ II.3 Error Analysis
We assume the transformation $\mathbf{M}$ has a bounded norm and note that its norm is independent of the choice of model $\mathcal{M}$. The overall error certainly depends on $\bm{\theta}$ and its norm, but, as we show below, generic features of generalization are largely driven by the operator norms of the block components of the GAD, Eq. (30) rather than by specific choices of $\bm{\theta}$.
Since we are particularly interested in how these contributions depend on the number of model parameters, consider the situation where a given decomposition $\Theta=\mathcal{M}\oplus\mathcal{U}$ of parameter space is changed by moving one column $\bm{\varphi}$ out of the nescience matrix and into the design matrix. We introduce the notation $\mathcal{M}(m)$ in which we make explicit the dependence on the particular decomposition. With this convention, we now analyze each of the elements of the GAD in turn.
II.3.1 Data Insufficiency
Data insufficiency refers to the upper left block of Eq. (30), $\mathbf{I}_{m}-\mathbf{B}$. It is straightforward (see Section V.3 for details) to show that this is operator is the orthogonal projection of $\mathcal{M}$ onto the null space $\mathcal{N}\subseteq\mathcal{M}$ of $\mathbf{M}_{\mathcal{T}\mathcal{M}}$. To simplify notation, we denote this projection operator $\mathbf{P}_{\mathcal{N}}$. As a projection operator, its induced operator norm is always one.
Notice that $\|\mathbf{P}_{\mathcal{N}}\bm{\theta}_{\mathcal{M}}\|^{2}=\|\bm{\theta}_{\mathcal{M}}-\hat{\bm{\theta}}_{\mathcal{M}}\|^{2}$ is the squared bias in $\bm{\theta}_{\mathcal{M}}$. Let $\mathcal{N}(m)$ denote the null space of $MTM(m)$ for a particular model decomposition. Since $\mathcal{N}(m)\subseteq\mathcal{N}(m+1)$, it follows that the contribution $\|\mathbf{P}_{\mathcal{N}}\bm{\theta}_{\mathcal{M}}\|$ from parameter bias is a nondecreasing function of $m$.
In most generic cases, $\mathbf{M}_{\mathcal{T}\mathcal{M}}(m)$ has full column rank up to the interpolation threshold, $m=n$. In this case, data insufficiency is exactly zero for $m\leq n$, as depicted in Figure 2. This is a restatement of the well-known fact that the linear least squares estimator is generically unbiased in the classical regime. Data insuffiency quantifies the bias in the parameter estimates. In the current formulation, this bias originates when the Moore-Penrose psuedo-inverse is not a left inverse of $\mathbf{M}_{\mathcal{T}\mathcal{M}}$, which is generic when $\mathbf{M}_{\mathcal{T}\mathcal{M}}$ is wide and the model is over-parameterized. More generally, however, whenever $\mathbf{M}_{\mathcal{T}\mathcal{M}}$ is not full-column rank, the Moore-Penrose pseudoinverse leads to parameter bias and data insufficiency. In other scenarios, the learning algorithm can similarly introduce parameter bias and data insufficiency.
II.3.2 Parameter Insufficiency
Parameter insufficiency refers to the lower right block of Eq. (30). It is the most straightforward to analyze as it is simply the identity operator on the nescient parameters, $\mathbf{I}_{\mathcal{U}}$. Except in the trivial and uninteresting case that $\dim\mathcal{U}=0$, its operator norm is one. The contribution to the error from the parameter insufficiency is simply the norm of the nescient parameters, $\|\bm{\theta}_{\mathcal{U}}\|^{2}$. Parameter insufficiency dominates for small values of $m$, reflecting the fact most of the signal is nescient and the model lacks the capacity to capture if faithfully. It follows that parameter insufficiency is a non-increasing function of $m$ since it decreases by exactly $|\theta_{m+1}|^{2}$ at each step.
II.3.3 Generalized Aliasing
Finally, we consider contributions from the generalized aliasing operator $\mathbf{A}$. This is the most complicated contribution, but is the source of non-trivial generalization curves such as double or multiple descent, or multiple peaks from structured data.